OpenStudy (anonymous):

Evaluate the given integral using the substitution (or method) indicated. e^(x/8) dx; u = x/8

OpenStudy (anonymous):

u = x/8 du = 1/8 So the integral e^(x/8) = 1/8 e^u du

OpenStudy (anonymous):

\[\int\limits_{}^{}e^{\frac{ x }{ 8 }}dx\] Sound right?

OpenStudy (anonymous):

= 8 e^(x/8) + C

OpenStudy (anonymous):

ASSUME t=x/8 the du=dx/8 substitute dx=8du